OpenStudy (anonymous):

What is the slope-intercept form of this equation? 3x + 9y = 18 @Alexandra675

OpenStudy (btaylor):

To put it into slope intercept form means solving for y. so, how would you go about that?

OpenStudy (anonymous):

\[y=2-\frac{ x }{ 3 }\]

OpenStudy (anonymous):

y= 1/3x +2?

OpenStudy (btaylor):

@eflatt13 @Alexandra675 is right. Remember that you are subtracting the "x" term from both sides.

OpenStudy (btaylor):

But in terms of precise slope-intercept form, your form is correct (y=mx+b). so, I'd write it \(y=-\frac{x}{3} + 2\) and it's perfect

OpenStudy (mathstudent55):

First, divide both sides by 3 since every term is divisible by 3. \(3x + 9y = 18\) \(x + 3y = 6\) Now subtract x from both sides: \(3y = -x + 6\) Now divide both sides by 3 to get y by itself: \(y = -\dfrac{1}{3}x + 2\) Now compare the answer with your choices and choose the correct one.